OpenStudy (anonymous):

What is the solution set for |3x + 12| + 5 = 23?

OpenStudy (anonymous):

|3x+12|=18 3x+12=18 or 3x+12=-18 3x=6 x=2 or 3x=-30 x=-10 do you understand?

OpenStudy (anonymous):

Yes thankyou

OpenStudy (anonymous):

welcome :)